The average age of teachers in a certain town is 34 with a standard deviation of
4. If the principal of ABC college employs 100 teachers, find the probability that
the average age of these teachers is less than 35
1
Expert's answer
2021-07-18T17:37:02-0400

"\mu=34 \\

\sigma = 4 \\

n=100 \\

P(\bar{x}<35) = P(Z< \frac{35-34}{4 / \sqrt{100}}) \\

= P(Z< 2.5) \\

= 0.9937 \\

= 99.37 \; \%"